OpenStudy (anonymous):

I'LL GIVE A MEDAL!!! PLEASE HELPP >:( Judah and 4 of his friends shared some apples. Each of the five friends had one and one-eighth of an apple. How many apples did they eat? four and four-ninths five and five-eighths five and one-eighth six and one-eighth

OpenStudy (anonymous):

5 (people) 1 1/8 (apples) \[5 \times 1\frac{ 1 }{ 8 }\]\[(5 \times 1) + (5 \times \frac{ 1 }{ 8 })\] Got it from here?

OpenStudy (anonymous):

No lol

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

Multiply then add...

OpenStudy (anonymous):

Is it

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

YES! B. five and five-eighths

OpenStudy (anonymous):

LOL